OpenStudy (anonymous):

Can anyone help me with this problem? "You are traveling in a car toward a certain mountain at a speed of 80 km/hr. Your viewing angle to the mountain is 0.5 degrees. Fifteen minutes later, the same mountain has a viewing angle of 1 degree. About how tall is the mountain?

OpenStudy (anonymous):

okay did you start this problem

OpenStudy (anonymous):

I tried, but I don't think I am doing this right. All I have is that the ship traveled 20 km in 15 minutes.

OpenStudy (anonymous):

was there some picture that came with this problem? I am having trouble understanding what 'viewing angle' means.

OpenStudy (anonymous):

Well I had to draw this diagram. It's a basic right triangle with the mountain as perpendicular to the ground and the distance away as a leg. Then there is a hypotneuse. The viewing angle is the number of degrees above the horizontal. Sorry I am terrible at explaining but I hope you understand what I did.

OpenStudy (anonymous):

I got it thanks

OpenStudy (anonymous):

okay :) thanks

OpenStudy (anonymous):

Have you learn law of cosine and law of sine yet?

OpenStudy (anonymous):

Yeah sin theta = y/r and cos theta = x/r right?

OpenStudy (anonymous):

no that's just sine and cos(only used in right triangle) Law of sine and cosine can be used in any triangle http://en.wikipedia.org/wiki/Law_of_cosines

OpenStudy (anonymous):

oh okay I get it